# A Survey of Large Language Model Empowered Agents for Recommendation and Search: Towards Next-Generation Information Retrieval

### v1

@@ 4.1.3. Representation Optimization
Specifically, AgentCF(Zhang et al., 2024e) treats both users and items as agents, each with its memory module to maintain the preferences and tastes of potential adopters. In addition, the study introduces a new collaborative learning method for simultaneously optimizing user agents and item agents and designs a collaborative reflection mechanism that enables agents to adjust their memories based on differences in real-world interaction records to more accurately reflect user behavior. KGLA converts the path information in the knowledge graph into natural language descriptions, enhances the language agent’s understanding of user preferences, and dynamically updates user memory based on the interaction between users and items during the simulation phase. At the same time, it uses information such as brands, categories, and product features in the knowledge graph to generate detailed item descriptions, and more accurately characterizes item features by analyzing the path relationship between users and items, thereby improving the personalization and accuracy of the recommendation. In the Rec4Agentverse (Zhang et al., 2024a) framework, items are converted into interactive, intelligent, and proactive LLM agents that can dynamically acquire user preferences and continuously update their own characteristics through multiple rounds of dialogue. Unlike traditional static items, LLM agents can provide multi-dimensional knowledge representation and enhance service content through collaboration with other agents. For example, a travel agent can not only recommend itineraries based on user interests but also cooperate with other agents to provide more comprehensive services. In addition, agent items can accurately model user preferences and improve the accuracy of personalized recommendations and user experience.
@@ 4.1.4. System Integration
Traditional RSs often use a batch processing architecture to regularly update models and recommendation results. For example, models can be retrained every day or every week to reflect the latest user preferences. Data cleaning, feature engineering, model selection, and so on need to be considered separately during system integration. In contrast, LLM agents significantly reduce the need for manual intervention by learning and dynamically adjusting recommendation strategies in real time and supporting an instant feedback mechanism, which not only improves the efficiency of the recommendation system but also enhances the user experience. In addition, LLM agents also support a team collaboration framework where each agent can focus on different data sources or tasks. Through multi-level and multi-dimensional collaboration and information sharing between these agents, the system can more comprehensively understand user preferences and provide highly personalized and context-related recommendation content. From the perspective of framework structure, these studies can be divided into single-agent work (Wang et al., 2023a; Huang et al., 2023a; Jin et al., 2023; Shi et al., 2024), multi-agent collaboration (Fang et al., 2024; Wang et al., 2024d; Nie et al., 2024).
@@ 4.1.5. Environment Simulation
Specifically, iEvaLM (Wang et al., 2023b) represents an innovative interactive evaluation framework leveraging Large Language Models (LLMs), which incorporates an LLM-based user simulator to comprehensively model diverse system-user interaction scenarios. Extensive experimental evaluations conducted on two publicly available Conversational Recommender System (CRS) datasets demonstrate that iEvaLM achieves substantial performance enhancements over conventional evaluation methodologies. Notably, this framework introduces a novel dimension by incorporating the assessment of recommendation interpretability, thereby addressing a critical aspect often overlooked in traditional evaluation paradigms. Agent4Rec (Zhang et al., 2024b) generates 1,000 LLM-enabled agents with different social traits and preferences by initializing from the MovieLens-1M dataset. The agents interact with the RS on a page-by-page basis, performing operations such as watching, rating, evaluating, exiting, and interviewing. Experimental results show that the generative agent can effectively identify and respond to items that match user preferences, and the feedback provided by the agent can be used as enhanced data for iterative training and improvement of the recommendation strategy, thus forming a dual-track approach for comprehensive evaluation of recommendation algorithms. Yoon et al. (Yoon et al., 2024) introduced a novel evaluation framework designed to quantitatively assess the behavioral alignment between LLMs and human users in CRS. This comprehensive protocol encompasses five distinct evaluation tasks: (1) item selection for discussion, (2) binary preference expression, (3) open-ended preference articulation, (4) recommendation solicitation, and (5) feedback provision. Through systematic experimentation, their study not only identifies significant behavioral discrepancies between LLM-generated responses and human interactions but also provides empirical insights into mitigating these deviations through optimized model selection criteria and advanced prompting techniques. Zhu et al. (Zhu et al., 2024a) conducted an analysis of the inherent limitations in current LLM-based user simulation approaches, particularly focusing on the issue of inadvertent data leakage. To address these challenges, they proposed SimpleUserSim, an innovative framework that implements a controlled information disclosure mechanism. This novel simulator employs a direct conversational strategy to steer dialogue topics toward target items while maintaining strict information constraints - specifically, it restricts access to target item titles and operates solely on item attribute information until successful recommendation completion. The RecAgent (Wang et al., 2023d) framework has a detailed definition of LLM agent simulation users, including the Profile Module (ensuring that the agent has a unique personality), Memory Module (remember past behaviors and evolve dynamically in the environment), and Action Module (including RS-related searches, browsing, clicks, and page turns, as well as some social behaviors), so that it can effectively simulate real human behavior. The CSHI framework (Zhu et al., 2024b) introduces control mechanisms, enhances the scalability of simulation, and incorporates elements of human-computer collaboration, enabling agents to flexibly adjust their behavior based on real-time context and historical data, thereby providing a more accurate personalized experience in conversational RSs. Beyond acting as user simulators in conversational recommendation scenarios, LLM agents have further broadened their application areas, playing a vital role in reinforcement learning and attacks on RS. In the latest research, the SUBER framework (Corecco et al., [n. d.]) ingeniously utilizes LLM agents to simulate human behavior, thereby constructing a synthetic environment specifically for the training and evaluation of reinforcement learning-based RSs. This innovative approach effectively mitigates the high training costs associated with the scarcity of online data and addresses the challenges in evaluating RS models, particularly in accurately measuring model performance without directly impacting the real user experience. Zhang et al. (Zhang et al., 2024f) addresses the shortcomings of current user simulators such as SUBER and Agent4Rec (including high computational cost, susceptibility to ”hallucinations”, and failure to fully capture the complex user-system interaction dynamics) by explicitly modeling user preferences, combining the advantages of logical reasoning and statistical learning, and designing an effective evaluation framework to improve the reliability and efficiency of the simulators. CheatAgent (Ning et al., 2024) first analyzes the parts of the text that have the most significant impact on recommendation results. Then, it utilizes an LLM agent to generate adversarial perturbations that can mislead the RS and optimize the attack strategy through prompt tuning. Finally, it fine-tunes the LLM’s attack strategy based on feedback from the target system to enhance the attack’s effectiveness.
@@ 4.2.1. Taxonomy Introduction
Typically, the interaction between users and search engines can be divided into four key steps: goal, query writing, search execution, and result understanding (White, 2024). Figure 4 shows the five domains used by LLM agents for search tasks, which can correspond to the typical IR interaction process mentioned 2. Detailed descriptions of each domain are as follows:
@@
  • Query Rewriters: LLM agents are good at refining user queries iteratively, making them clearer and more specific, leading to more accurate search results. In this case, LLM agents serve as the query modeling modules.
@@
  • Results Synthesizers: LLM agents are suitable for summarizing large amounts of text and helping users quickly grasp the essential details to make decisions. In this case, LLM agents bring new abilities into the traditional searching process.
@@
We categorize research work related to search tasks according to the descriptions in the abovementioned domains and list typical works in the five domains in Table 2.
@@ 4.2.2. Task Decomposers
The existing search technologies typically focus on learning the relevance between user queries and documents without any thoughtful processing. This limitation is critical for complex search tasks. For example, when a user searches for “where to go on vacation?”, his/her expectation might be a comparison of various travel destinations, details on travel itinerary, and hotel/flight prices. As discussed in Section 3, LLM agents possess strong reasoning and task decomposition abilities, enabling them to collect and aggregate information step by step, providing comprehensive results to the user.
@@
Laser (Ma et al., 2023b) introduces a method that views web navigation tasks as state-space exploration. LLM agents decompose tasks into the pre-defined set of states to handle unfamiliar scenarios and flexible backtracking. Deng et al. (Deng et al., 2024c) propose a new task called Conversational Web Navigation, which requires complex multi-turn interactions with both users and the environment. In cases where tasks involve intricate actions, the agents will initially decompose the task into smaller sub-tasks. The decomposition results will later be verified and corrected by human annotators. Gur et al. (Gur et al., 2023a) introduce a LLM agent named WebAgent. WebAgent has the capability to decompose instructions into standardized sub-instructions and plan ahead, which enables it to better understand and execute complex tasks. For instance, in a complex web operation task, it can first identify the main steps that need to be executed and then carry out these steps sequentially. SteP (Sodhi et al., 2023) proposes a policy stack to better coordinate different actions and accomplish complex tasks. By decomposing tasks into various policy layers, each policy can focus on specific aspects of the task, thereby improving efficiency and accuracy in task execution. For example, in handling web tasks, one policy might be responsible for page navigation while another policy might handle form filling. Together, these policies form a policy stack that collaboratively completes the entire task. Koh et al. (Koh et al., 2024b) propose an inference-time search algorithm where agents do a tree search progress to decide which action to take according to the current state. This method helps LLM agents better perform multi-step planning in interactive web environments. Knowagent (Zhu et al., 2024c) utilizes an action knowledge base and a knowledge-driven self-learning approach to guide the action paths during the planning phase. The action knowledge base furnishes the agent with extensive action knowledge, allowing for more rational action selection during planning. Meanwhile, the knowledge-driven self-learning approach enables the agent to continuously learn and refine its planning strategies, enhancing overall planning performance. Xie et al. (Xie and Zou, 2024) introduce a human-like reasoning framework to improve the planning ability that can help LLM agents when solving multi-phase problems like travel planning. Agent Q (Putta et al., 2024) proposes a framework that integrates guided Monte Carlo Tree Search (MCTS) with iterative fine-tuning for agents. This approach enables LLM agents to learn effectively from all trajectories, including both successful and unsuccessful ones, thereby enhancing their ability to generalize in complex, multi-step reasoning tasks.
@@ 4.2.3. Query Rewriters
In traditional search scenarios, users must decide on the search terms themselves and master complex syntax to emphasize which keywords should necessarily be included or excluded from the search, and the length and context of these queries are usually limited. These limitations can prevent users from fully expressing their true intentions. However, these obstacles can be solved through LLM agents. LLM agents allow users to input longer context and fully express their intentions and then formulate more effective queries by themselves, thereby enhancing search efficiency.
CoSearchAgent (Gong et al., 2024) is a lightweight collaborative search agent powered by LLMs and serves as a search plugin on the instant message platform Slack. This agent can autonomously extract the user’s search intent based on the conversation between users, formulate it into a query, and then invoke a search engine to return the results. The entire process does not require the user to explicitly input a query, thereby avoiding disruption to the conversational experience. Joko et al. (Joko et al., 2024) propose a scheme called LAPS (LLM-Augmented Personalized Self-Dialogue) to gather personalized dialogues, which can be used to train preference extraction and personalized response generation. As more conversational data is collected, LLM agents can gradually learn about users’ various preferences, for instance, a user can be vegetarianism. Subsequently, when the same user inquires about recipes, the agents can account for vegetarian needs and directly search for vegetarian-related rather than normal recipes. Aliannejadi et al. (Aliannejadi et al., 2024) organize a competition on how to utilize users’ prior interactions and present context and observed two main pipelines, namely, multi-turn retrieve-then-generate (R→G) and generate-retrieve-then-generate (G→R→G). In this context, ”G” includes conversational query expansion and conversational rewriting. Jain et al. (Jain et al., 2024) propose an approach utilizing Retrieval Augmented Generation (RAG) to enable LLM agents to enrich user queries with additional information. Specifically, by leveraging RAG, the agents can augment user queries with pertinent details sourced from GitHub repositories.
@@ 4.2.4. Action Executors
In summary, recent research has focused on two key areas. Firstly, how to equip LLMs agents with a broader range of tools, such as code execution and web retrieval. Secondly, how to make LLM agents better select the most appropriate tools from various options.
@@ 4.2.5. Results Synthesizers
PersonaRAG (Zerhoudi and Granitzer, 2024) is an innovative framework that incorporates user-centric agents to select retrieval and generation based on real user preference. This framework works through three steps: retrieval, user interaction analysis, and cognitive dynamic adaptation. PersonaRAG outperforms baseline models, delivering personalized answers that better address user needs on various question-answering datasets. ChatCite (Li et al., 2024a) is an LLM agent guided by human workflows for comparative literature summarization. The agent mimics the human workflow to extract key elements from relevant literature and generate summaries through a reflective incremental mechanism. This approach enables ChatCite to achieve a deeper understanding of the literature content and produce more targeted and comparative summaries. PaSa (He et al., 2025) introduces a hierarchical architecture comprising Crawler and Selector agents to address complex academic search challenges. The Crawler agent dynamically generates search queries and expands citation networks through tool invocation, while the Selector agent evaluates paper relevance via fine-grained scoring. To optimize this process, a novel conversational PPO algorithm is developed to handle sparse rewards in long-horizon search tasks. The system is trained on AutoScholarQuery, a synthetic dataset with 35k AI-domain queries, and validated on RealScholarQuery containing real-world research needs. Experiments show that PaSa-7B achieves 37.78% higher recall@20 than GPT-4o-enhanced Google Search, demonstrating superior capability in resolving specialized academic queries.

### v2

@@ 4.1.3. Representation Optimization
Specifically, AgentCF(Zhang et al., 2024e) treats both users and items as agents, each with its memory module to maintain the preferences and tastes of potential adopters. In addition, the study introduces a new collaborative learning method for simultaneously optimizing user agents and item agents and designs a collaborative reflection mechanism that enables agents to adjust their memories based on differences in real-world interaction records to more accurately reflect user behavior. KGLA (Guo et al., 2024) converts the path information in the knowledge graph into natural language descriptions, enhances the language agent’s understanding of user preferences, and dynamically updates user memory based on the interaction between users and items during the simulation phase. At the same time, it uses information such as brands, categories, and product features in the knowledge graph to generate detailed item descriptions, and more accurately characterizes item features by analyzing the path relationship between users and items, thereby improving the personalization and accuracy of the recommendation. In the Rec4Agentverse (Zhang et al., 2024a) framework, items are converted into interactive, intelligent, and proactive LLM agents that can dynamically acquire user preferences and continuously update their own characteristics through multiple rounds of dialogue. Unlike traditional static items, LLM agents can provide multi-dimensional knowledge representation and enhance service content through collaboration with other agents. For example, a travel agent can not only recommend itineraries based on user interests but also cooperate with other agents to provide more comprehensive services. In addition, agent items can accurately model user preferences and improve the accuracy of personalized recommendations and user experience.
@@ 4.1.4. System Integration
Traditional RSs often use a batch processing architecture to regularly update models and recommendation results. For example, models can be retrained every day or every week to reflect the latest user preferences. Data cleaning, feature engineering, model selection, and so on need to be considered separately during system integration. In contrast, LLM agents significantly reduce the need for manual intervention by learning and dynamically adjusting recommendation strategies in real-time and supporting an instant feedback mechanism, which not only improves the efficiency of the recommendation system but also enhances the user experience. In addition, LLM agents also support a team collaboration framework where each agent can focus on different data sources or tasks. Through multi-level and multi-dimensional collaboration and information sharing between these agents, the system can more comprehensively understand user preferences and provide highly personalized and context-related recommendation content. From the perspective of framework structure, these studies can be divided into single-agent work (Wang et al., 2023a; Huang et al., 2023a; Jin et al., 2023; Shi et al., 2024), multi-agent collaboration (Fang et al., 2024; Wang et al., 2024d; Nie et al., 2024).
@@ 4.1.5. Environment Simulation
Specifically, iEvaLM (Wang et al., 2023b) represents an innovative interactive evaluation framework leveraging LLMs, incorporating an LLM-based user simulator to comprehensively model diverse system-user interaction scenarios. Extensive experimental evaluations conducted on two publicly available Conversational Recommender System (CRS) datasets demonstrate that iEvaLM achieves substantial performance enhancements over conventional evaluation methodologies. Notably, this framework introduces a novel dimension by incorporating the assessment of recommendation interpretability, thereby addressing a critical aspect often overlooked in traditional evaluation paradigms. Agent4Rec (Zhang et al., 2024b) generates 1,000 LLM-enabled agents with different social traits and preferences by initializing from the MovieLens-1M dataset. The agents interact with the RS on a page-by-page basis, performing operations such as watching, rating, evaluating, exiting, and interviewing. Experimental results show that the generative agent can effectively identify and respond to items that match user preferences, and the feedback provided by the agent can be used as enhanced data for iterative training and improvement of the recommendation strategy, thus forming a dual-track approach for comprehensive evaluation of recommendation algorithms. Yoon et al. (Yoon et al., 2024) introduced a novel evaluation framework designed to quantitatively assess the behavioral alignment between LLMs and human users in CRS. This comprehensive protocol encompasses five distinct evaluation tasks: (1) item selection for discussion, (2) binary preference expression, (3) open-ended preference articulation, (4) recommendation solicitation, and (5) feedback provision. Through systematic experimentation, their study not only identifies significant behavioral discrepancies between LLM-generated responses and human interactions but also provides empirical insights into mitigating these deviations through optimized model selection criteria and advanced prompting techniques. Zhu et al. (Zhu et al., 2024a) conducted an analysis of the inherent limitations in current LLM-based user simulation approaches, particularly focusing on the issue of inadvertent data leakage. To address these challenges, they proposed SimpleUserSim, an innovative framework that implements a controlled information disclosure mechanism. This novel simulator employs a direct conversational strategy to steer dialogue topics toward target items while maintaining strict information constraints - specifically, it restricts access to target item titles and operates solely on item attribute information until successful recommendation completion. The RecAgent (Wang et al., 2023d) framework has a detailed definition of LLM agent simulation users, including the Profile Module (ensuring that the agent has a unique personality), Memory Module (remember past behaviors and evolve dynamically in the environment), and Action Module (including RS-related searches, browsing, clicks, and page turns, as well as some social behaviors), so that it can effectively simulate real human behavior. The CSHI framework (Zhu et al., 2024b) introduces control mechanisms, enhances the scalability of simulation, and incorporates elements of human-computer collaboration, enabling agents to flexibly adjust their behavior based on real-time context and historical data, thereby providing a more accurate personalized experience in conversational RSs. Beyond acting as user simulators in conversational recommendation scenarios, LLM agents have further broadened their application areas, playing a vital role in reinforcement learning and attacks on RS. In the latest research, the SUBER framework (Corecco et al., [n. d.]) ingeniously utilizes LLM agents to simulate human behavior, thereby constructing a synthetic environment specifically for the training and evaluation of reinforcement learning-based RSs. This innovative approach effectively mitigates the high training costs associated with the scarcity of online data and addresses the challenges in evaluating RS models, particularly in accurately measuring model performance without directly impacting the real user experience. Zhang et al. (Zhang et al., 2024f) addresses the shortcomings of current user simulators such as SUBER and Agent4Rec (including high computational cost, susceptibility to “hallucinations”, and failure to fully capture the complex user-system interaction dynamics) by explicitly modeling user preferences, combining the advantages of logical reasoning and statistical learning, and designing an effective evaluation framework to improve the reliability and efficiency of the simulators. CheatAgent (Ning et al., 2024) first analyzes the parts of the text that have the most significant impact on recommendation results. Then, it utilizes an LLM agent to generate adversarial perturbations that can mislead the RS and optimize the attack strategy through prompt tuning. Finally, it fine-tunes the LLM’s attack strategy based on feedback from the target system to enhance the attack’s effectiveness.
@@ 4.2.1. Taxonomy Introduction
Typically, the interaction between users and search engines can be divided into four key steps: user intent, query writing, search execution, and result understanding (White, 2024). Figure 4 shows the five domains utilizing LLM agents in search tasks, which can correspond to the typical IR interaction process mentioned 2. Detailed descriptions of each domain are as follows:
@@
  • Query Rewriters: LLM agents are good at refining user queries to make them clearer and more specific, leading to more accurate search results. In this case, LLM agents serve as the query modeling modules.
@@
  • Results Synthesizers: LLM agents are suitable for summarizing large amounts of search results and helping users quickly grasp the essential details to make decisions. In this case, LLM agents bring new abilities into the traditional searching process.
@@
We categorize research work related to search tasks according to the descriptions in the domains mentioned above and list typical works in the five domains in Table 2.
@@ 4.2.2. Task Decomposers
The existing search technologies typically focus on learning the relevance between user queries and documents without any thoughtful processing. This limitation is critical for complex search tasks. For example, when a user searches for “where to go on vacation?”, his/her expectation might be a comparison of various travel destinations, details on travel itineraries, and hotel/flight prices. As discussed in Section 3, LLM agents possess strong reasoning and task decomposition abilities, enabling them to collect and aggregate information step by step and provide comprehensive results to the user.
@@
Laser (Ma et al., 2023b) introduces a method that views web navigation tasks as state-space exploration. LLM agents decompose tasks into a predefined set of states to handle unfamiliar scenarios and flexible backtracking. Deng et al. (Deng et al., 2024c) propose a new task called Conversational Web Navigation, which requires complex multi-turn interactions with both users and the environment. In cases where tasks involve intricate actions, the agents will initially decompose the task into smaller sub-tasks. The decomposition results will later be verified and corrected by human annotators. Gur et al. (Gur et al., 2023a) introduce a LLM agent named WebAgent. WebAgent has the capability to decompose instructions into standardized sub-instructions and plan ahead, which enables it to better understand and execute complex tasks. For instance, in a complex web operation task, it can first identify the main steps that need to be executed and then carry out these steps sequentially. SteP (Sodhi et al., 2023) proposes a policy stack to better coordinate different actions and accomplish complex tasks. By decomposing tasks into various policy layers, each policy can focus on specific aspects of the task, thereby improving efficiency and accuracy in task execution. For example, in handling web tasks, one policy might be responsible for page navigation, while another policy might handle form filling. Together, these policies form a policy stack that collaboratively completes the entire task. Koh et al. (Koh et al., 2024b) propose an inference-time search algorithm where agents do a tree search to decide which action to take according to the current state. This method helps LLM agents better perform multi-step planning in interactive web environments. Knowagent (Zhu et al., 2024c) utilizes an action knowledge base and a knowledge-driven self-learning approach to guide the action paths during the planning phase. The action knowledge base furnishes the agent with extensive action knowledge, allowing for more rational action selection during planning. Meanwhile, the knowledge-driven self-learning approach enables the agent to continuously learn and refine its planning strategies, enhancing overall planning performance. Xie et al. (Xie and Zou, 2024) introduce a human-like reasoning framework to improve the planning ability that can help LLM agents when solving multi-phase problems like travel planning. Agent Q (Putta et al., 2024) proposes a framework that integrates guided Monte Carlo Tree Search (MCTS) with iterative fine-tuning for agents. This approach enables LLM agents to learn effectively from all trajectories, including both successful and unsuccessful ones, thereby enhancing their ability to generalize in complex, multi-step reasoning tasks.
@@ 4.2.3. Query Rewriters
In traditional search scenarios, users must decide on the search terms themselves and master complex syntax to emphasize which keywords should necessarily be included or excluded from the search, and the length and context of these queries are usually limited. These limitations can prevent users from fully expressing their true intentions. However, these obstacles can be solved through LLM agents. LLM agents allow users to input longer context to fully express their intentions and then formulate more effective queries by themselves, thereby enhancing search efficiency.
CoSearchAgent (Gong et al., 2024) is a lightweight collaborative search agent powered by LLMs and serves as a search plugin on the instant message platform Slack. This agent can autonomously extract the user’s search intent based on the conversation between users, formulate it into a query, and then invoke a search engine to return the results. The entire process does not require the user to explicitly input a query, thereby avoiding disruption to the conversational experience. Joko et al. (Joko et al., 2024) propose a scheme called LAPS (LLM-Augmented Personalized Self-Dialogue) to gather personalized dialogues, which can be used to train preference extraction and personalized response generation. As more conversational data is collected, LLM agents can gradually learn about users’ various preferences. For instance, when a vegetarian user inquires about recipes, the agents can account for vegetarian needs and directly search for vegetarian-related rather than normal recipes. Aliannejadi et al. (Aliannejadi et al., 2024) organize a competition on how to utilize users’ prior interactions and present context and observed two main pipelines, namely, multi-turn retrieve-then-generate (R→G) and generate-retrieve-then-generate (G→R→G). In this context, “G” includes conversational query expansion and conversational rewriting. Jain et al. (Jain et al., 2024) propose an approach utilizing Retrieval Augmented Generation (RAG) to enable LLM agents to enrich user queries with additional information. Specifically, by leveraging RAG, the agents can augment user queries with pertinent details sourced from GitHub repositories.
@@ 4.2.4. Action Executors
In summary, recent researches have focused on two key areas. Firstly, how to equip LLMs agents with a broader range of tools, such as code execution and web retrieval. Secondly, how to make LLM agents better select the most appropriate tools from various options.
@@ 4.2.5. Results Synthesizers
PersonaRAG (Zerhoudi and Granitzer, 2024) is an innovative framework that incorporates user-centric agents to select retrieval and generation based on real user preference. This framework works through three steps: retrieval, user interaction analysis, and cognitive dynamic adaptation. PersonaRAG outperforms baseline models, delivering personalized answers that better address user needs on various question-answering datasets. ChatCite (Li et al., 2024a) is an LLM agent guided by human workflows for comparative literature summarization. The agent mimics the human workflow to extract key elements from relevant literature and generate summaries through a reflective incremental mechanism. This approach enables ChatCite to achieve a deeper understanding of the literature content and produce more targeted and comparative summaries. PaSa (He et al., 2025) introduces a hierarchical architecture comprising Crawler and Selector agents to address complex academic search challenges. The Crawler agent dynamically generates search queries and expands citation networks through tool invocation, while the Selector agent evaluates paper relevance via fine-grained scoring. To optimize this process, a novel conversational PPO algorithm is developed to handle sparse rewards in long-horizon search tasks.